# Detecting Out-of-distribution through the Lens of Neural Collapse

### v4

@@ 1 Introduction
In this work, we focus on post-hoc methods, which address OOD detection independently of the training process. One line of prior work designs OOD scores over model output space (Djurisic et al., 2022; Hendrycks et al., 2019; Liang et al., 2018; Liu et al., 2020; Sun et al., 2021; Sun and Li, 2022) and another line of work focuses on the feature space, where OOD samples are observed to deviate from the clusters of ID samples (Lee et al., 2018; Mahalanobis, 2018; Sun et al., 2022; Tack et al., 2020). While existing research has made strides in OOD detection, they still face two major challenges: 1) maintaining detection effectiveness across different scenarios, and 2) ensuring computational efficiency for real-world deployment. For example, both output space and feature space methods suffer from performance discrepancy across different classification tasks, as shown in Table 1(a). Specifically, state-of-art algorithms on CIFAR-10 (Krizhevsky et al., 2009) OOD benchmarks perform suboptimally on ImageNet (Deng et al., 2009) OOD benchmarks, and vice versa. Such discrepancy is also observed across different architectures, as shown in Table 2. In addition, feature space methods, which rely on auxiliary models, raise efficiency concerns. For example, Lee et al. (2018) learns a Gaussian mixture model from training features and detects OOD based on Mahalanobis distance Mahalanobis (2018); Sun et al. (2022) records the training features and measures OOD-ness based on the k-th nearest neighbor distance to the training features. As shown in Liu and Qin (2023), such reliance on auxiliary models introduces additional cost, posing challenges for time-critical applications.
@@
To this end, we aim to develop an efficient and versatile OOD detector. Specifically, we delve into the penultimate layer, i.e., the layer before the linear classification head, and revisit the observation that ID features tend to form clusters while OOD features reside apart. While this observation is well-established in prior literature Lee et al. (2018); Sun et al. (2022); Tack et al. (2020), the underlying mechanism remains largely unexplained. To better understand and characterize the observation, we take insights from Neural Collapse (Papyan et al., 2020). Neural Collapse characterizes the interplay between the linear classification head and the penultimate layer features as training epochs go to infinity. Notably, the Neural Collapse phenomenons are observed to be prevalent across canonical classification tasks under diverse architectures (see Appendix E). One particular phenomenon that Neural Collapse reveals is that features of each class gradually converge towards a single point during training. This explains the prevalent clustering observed on off-the-shelf models as a precursor to the final convergence. Therefore, we leverage the landscape of Neural Collapse to understand:
@@
Such a landscape of Neural Collapse sheds light on the geometric structure of ID clusters. Specifically, for ID test samples, drawn from the same distribution as training samples, we anticipate a similar clustering behavior towards the weight vectors and towards an ETF. Conversely, OOD samples do not undergo the same training process, which enables the model to align features with weight vectors and to expand features to accommodate the spatial structure of ETF in Neural Collapse. Therefore, we do not expect the model to effectively align the weight vectors learned from ID features with unseen OOD features. Nor do we anticipate the model to posit OOD features far from the origin to structure an ETF. To validate our hypotheses, we trace a model’s training stages in Figure 2. We observe that ID samples consistently cluster closer to the weight vectors than OOD samples. This observation is reinforced in our UMAP McInnes et al. (2018) visualization on an off-the-shelf CIFAR-10 classifier with ResNet-18 backbone in Figure 1 Left. Here, ID features cluster near predicted class weight vectors (marked by stars), while OOD features are distant. Combining our observation with the results in (Zhu et al., 2021), which show that the weight vectors form an ETF in training, we conclude that ID features are driven to structure the ETF during training, whereas OOD features have no incentive to expand in space to form an ETF. Note that the lack of incentive for OOD features to expand in space explains the well-established observation Huang et al. (2021); Park et al. (2023); Sun et al. (2022); Tack et al. (2020) that OOD features tend to reside closer to the origin.
@@
Notably, prior methods, e.g., KNN Sun et al. (2022), focus on ID clustering but do not explicitly consider feature distance to the origin. Such approaches fall short in scenarios like ImageNet benchmarks but yield superior performance in CIFAR-10 benchmarks in Table 1(a). Conversely, methods such as Energy Liu et al. (2020), Energy-based ASH Djurisic et al. (2022), and, Energy-based Scale Xu et al. (2023) inherently utilize feature distance to the origin by considering log-sum-exp of logits, yet largely overlook ID clustering. These approaches excel in some scenarios, e.g., ImageNet, but perform sub-optimally in others, e.g., CIFAR-10 (see Table 1(a)). Through the lens of Neural Collapse, we explain, connect, and complete prior methods under a holistic view, leading to improved efficiency and generalizability validated across diverse experiments.
@@
  • Experimental Analysis: We evaluate $\mathtt{NCI}$ across diverse classification tasks (CIFAR-10, CIFAR-100, ImageNet) and model architectures (ResNet, DenseNet Huang et al. (2017), ViT Dosovitskiy et al. (2020), Swin Liu et al. (2022)). Our $\mathtt{NCI}$ practically matches the latency of vanilla softmax-confidence detector while largely mitigating the performance discrepancy of existing methods across different classification tasks and model architectures.
@@ 3 OOD Detection through the Lens of Neural Collapse
In this section, we re-examine the observation in Lee et al. (2018); Sun et al. (2022) that ID features tend to form clusters while OOD features deviate from the clusters. We understand the phenomenon as a precursor to the Neural Collapse (Papyan et al., 2020) convergence. And leveraging the landscape revealed in Neural Collapse, we examine:
@@ 3.1 Neural Collapse: Convergence of Training Features
We first remark on (NC2) that an ETF achieves the maximum separation possible for globally centered equiangular vectors Papyan et al. (2020) and extends in space, as visualized in Figure 1 Right. Since training features converge towards an ETF, they need to have sufficient norms to accommodate the spatial arrangement.
@@ 3.2 Geometric Structure of the Clusters of ID Features
Additionally, we combine our observation with the results from (Zhu et al., 2021), showing that the weight vectors form an ETF during training. Our observed proximity to the weight vectors thus also validates the clustering of ID features near an ETF and the divergence of OOD from this structure. The lack of structure and incentives to extend in space explains the relatively smaller norm of OOD features.
@@ 3.3 Out-of-Distribution Detection
While $\mathtt{pScore}$ enhances efficiency and effectiveness, its performance is intrinsically contingent on the strength of ID clustering. Such contingency, widely exhibited by clustering-based methods Lee et al. (2018); Sun et al. (2022); Tack et al. (2020), poses challenges on classifiers with less pronounced ID clustering, such as ImageNet ResNet-50 in Section 4.1. To mitigate such discrepancy, we complement $\mathtt{pScore}$ by considering the distance of ID clusters to the origin. Specifically, we enhance our proximity score by incorporating feature norms to filter out OOD near the origin, as illustrated in Figure 1 Right. Taking $\mathtt{L1}$ norm as an example, we define our detection score as $\mathtt{pScore}+\alpha\|\bm{h}\|_{1}$, where $\alpha$ controls the filtering strength and can be selected from a validation set as detailed in Section 4. We refer readers to Section 4.3 for the effect of different orders of $p$-norm. Thresholding on the detection score, we have Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$): A lower score indicates a higher chance of OOD-ness.
@@ 4 Experiments
(a) $\mathtt{NCI}$ remains effective while baselines exhibit discrepancy. $\uparrow$ and $\downarrow$ denotes better performance. Bold indicates the best results while underline denotes the 2nd and 3rd best results. Methods with * are hyperparameter-free. Scores, except for the most recent baselines – $\mathtt{fDBD}$, $\mathtt{NECO}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – are copied from OpenOOD report Zhang et al. (2023).
@@
In this section, we extensively demonstrate the versatility and efficiency of $\mathtt{NCI}$ across classification tasks: CIFAR-10, CIFAR-100 (see App. D), ImageNet, as well as model architectures: ResNet, DenseNet (see App. D), ViT, Swin. We compare $\mathtt{NCI}$ against thirteen baseline methods, including the most recent ones, and demonstrate that $\mathtt{NCI}$ effectively mitigates the generalization discrepancies of existing methods. Following the OpenOOD benchmark Zhang et al. (2023), we evaluate on six OOD sets for CIFAR-10 and CIFAR-100 classifiers and five for ImageNet classifiers. Performance is evaluated using two widely recognized metrics: the False Positive Rate at 95% True Positive Rate (FPR95) and the Area Under the Receiver Operating Characteristic Curve (AUROC). Lower FPR95 and higher AUROC values indicate better performance. We also report the per-image inference latency (in milliseconds) evaluated on a Tesla T4 GPU. In our experiments, other than the ablation study in Section 4.3, we use the $L1$-norm as the filtering term and select the filtering strength $\alpha$ from $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ based on a validation set generated per pixel from Gaussian $N(0,1)$, following Sun et al. (2021); Sun and Li (2022). In Section 4.3, we explore different choices of norm order and examine the effect of filtering on alternative clustering-based method. For detailed setups, please see Appendix A. We emphasize that our method and all baselines are post-hoc methods, while all models used are off-the-shelf and do not undergo prolonged training.
@@ 4.1 Versatility across Classification Tasks
We first assess the performance of $\mathtt{NCI}$ and baselines across CIFAR-10 and ImageNet classification tasks. The two tasks provide an ideal test bed for evaluating versatility, as they drastically differ in input resolution, number of classes, and classification accuracy. We use the ResNets from the OpenOOD benchmarkZhang et al. (2023): the CIFAR-10 classifier is a ResNet-18 with an accuracy of 95.06% and the ImageNet classifier is a ResNet-50 with an accuracy of 76.65%. Based on validation results, we set the filter strength $\alpha$ of the $L1$-norm to $10^{-2}$ for CIFAR-10 experiments and $10^{-3}$ for ImageNet experiments.
Datasets For CIFAR-10 experiments, We follow the OpenOOD split of CIFAR-10 test set and evaluate on the OpenOOD benchmarks, including CIFAR-100 Krizhevsky et al. (2009), Tiny ImageNet Le and Yang (2015), MNIST Deng (2012), SVHN Netzer et al. (2011), Texture (Cimpoi et al., 2014), and Places365 (Zhou et al., 2017). For ImageNet experiments, we follow the OpenOOD split of ImageNet test set and evaluate on the OpenOOD benchmarks, including SSB-hard Vaze et al. (2021), NINCO Bitterwolf et al. (2023), iNaturalist (Van Horn et al., 2018), Texture (Cimpoi et al., 2014), and OpenImage-O Wang et al. (2022).
Baselines In Table 1(a), we compare our method with thirteen baselines. Some baselines focus more on the CIFAR-10 Benchmark while others focus more focused on the Imagenet Benchmark. Based on performance, we categorize the baselines, besides the vanilla confidence-based $\mathtt{MSP}$ (Hendrycks and Gimpel, 2016), into two groups: the “CIFAR-10 Strong" baselines, including $\mathtt{ODIN}$ (Liang et al., 2018), $\mathtt{Energy}$ (Liu et al., 2020), $\mathtt{Mahalanobis}$ (Lee et al., 2018), $\mathtt{KNN}$(Sun et al., 2022), $\mathtt{ViM}$ (Wang et al., 2022), and $\mathtt{fDBD}$ Liu and Qin (2023); the “ImageNet Strong" baselines, including $\mathtt{GradNorm}$ (Huang et al., 2021), $\mathtt{NECO}$ Ammar et al. (2023), $\mathtt{React}$ (Sun et al., 2021), $\mathtt{Dice}$ (Sun and Li, 2022), $\mathtt{ASH}$ Djurisic et al. (2022), $\mathtt{Scale}$ Xu et al. (2023). For details of the baselines, please see Appendix C.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{NECO}$: Like our $\mathtt{NCI}$, $\mathtt{NECO}$ Ammar et al. (2023) is motivated by Neural Collapse. Similar to our $\mathtt{NCI}$ with filtering, $\mathtt{NECO}$ incorporates distance to the origin using max-logit. However, $\mathtt{NECO}$, exclusively focuses on the subspace analysis of features and does not utilize the classification head. Such subspace operations require expensive matrix multiplication, resulting in a noticeable inference latency shown in Table 1(b). Conversely, our $\mathtt{NCI}$ explores the interplay between features and the classification head, and thus integrates class-specific information revealed by Neural Collapse. Compared to $\mathtt{NECO}$, our $\mathtt{NCI}$ improves in both efficiency and effectiveness.
@@ 4.2 Versatility across Model Architectures
Now we assess the performance of $\mathtt{NCI}$ and baselines across different architectures. Particularly, we study two transformer-based models: ViT B/16 Dosovitskiy et al. (2020) and Swin-v2 Liu et al. (2022). Both transformers are finetuned on ImageNet, achieving an accuracy of 81.14% and 82.94% respectively. We follow the setup of the OpenOOD ImageNet Benchmark, detailed in Section 4.1. And based on validation results, we set the filter strength $\alpha$ of the $L1$ norm to $10^{-3}$ for both classifiers. In Table 2, we observe strong baselines on ViT – $\mathtt{KNN}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – exhibit discrepancy when applied to Swin v2, echoing the observations in Ammar et al. (2023). Conversely, our $\mathtt{NCI}$, even without filtering, improves baseline performance on Swin v2 while maintaining superior effectiveness on ViT. The addition of filtering further enhances overall generalizability. Our experiments on ResNet (Section 4.1) and DenseNet (Appendix D) further validate the versatility of $\mathtt{NCI}$ across model architectures.

### v5

@@ 1 Introduction
In this work, we focus on post-hoc methods, which address OOD detection independently of the training process. One line of prior work designs OOD scores over model output space (Djurisic et al., 2022; Hendrycks et al., 2019; Liang et al., 2018; Liu et al., 2020; Sun et al., 2021; Sun and Li, 2022) and another line of work focuses on the feature space, where OOD samples are observed to deviate from the clusters of ID samples (Lee et al., 2018; Mahalanobis, 2018; Sun et al., 2022; Tack et al., 2020). While existing research has made strides in OOD detection, they still face two major challenges: 1) maintaining detection effectiveness across different scenarios, and 2) ensuring computational efficiency for real-world deployment. For example, both output space and feature space methods suffer from performance discrepancy across different classification tasks, as shown in Table 1(a). Specifically, state-of-art algorithms on CIFAR-10 (Krizhevsky et al., 2009) OOD benchmarks perform suboptimally on ImageNet (Deng et al., 2009) OOD benchmarks, and vice versa. Such discrepancy is also observed across different architectures, as shown in Table 2. In addition, feature space methods, which rely on auxiliary models, raise efficiency concerns. For example, Lee et al. (2018) learns a Gaussian mixture model from training features and detects OOD based on Mahalanobis distance (Mahalanobis, 2018); Sun et al. (2022) records the training features and measures OOD-ness based on the k-th nearest neighbor distance to the training features. As shown in Liu and Qin (2023), such reliance on auxiliary models introduces additional cost, posing challenges for time-critical applications.
@@
To this end, we aim to develop an efficient and versatile OOD detector. Specifically, we delve into the penultimate layer, i.e., the layer before the linear classification head, and revisit the observation that ID features tend to form clusters while OOD features reside apart. While this observation is well-established in prior literature (Lee et al., 2018; Sun et al., 2022; Tack et al., 2020), the underlying mechanism remains largely unexplained. To better understand and characterize the observation, we take insights from Neural Collapse (Papyan et al., 2020). Neural Collapse characterizes the interplay between the linear classification head and the penultimate layer features as training epochs go to infinity. Notably, the Neural Collapse phenomenons are observed to be prevalent across canonical classification tasks under diverse architectures (see Appendix E). One particular phenomenon that Neural Collapse reveals is that features of each class gradually converge towards a single point during training. This explains the prevalent clustering observed on off-the-shelf models as a precursor to the final convergence. Therefore, we leverage the landscape of Neural Collapse to understand:
@@
Such a landscape of Neural Collapse sheds light on the geometric structure of ID clusters. Specifically, for ID test samples, drawn from the same distribution as training samples, we anticipate a similar clustering behavior towards the weight vectors and towards an ETF. Conversely, OOD samples do not undergo the same training process, which enables the model to align features with weight vectors and to expand features to accommodate the spatial structure of ETF in Neural Collapse. Therefore, we do not expect the model to effectively align the weight vectors learned from ID features with unseen OOD features. Nor do we anticipate the model to posit OOD features far from the origin to structure an ETF. To validate our hypotheses, we trace a model’s training stages in Figure 2. We observe that ID samples consistently cluster closer to the weight vectors than OOD samples. This observation is reinforced in our UMAP (McInnes et al., 2018) visualization on an off-the-shelf CIFAR-10 classifier with ResNet-18 backbone in Figure 1 Left. Here, ID features cluster near predicted class weight vectors (marked by stars), while OOD features are distant. Combining our observation with the results in Zhu et al. (2021), which show that the weight vectors form an ETF in training, we conclude that ID features are driven to structure the ETF during training, whereas OOD features have no incentive to expand in space to form an ETF. Note that the lack of incentive for OOD features to expand in space explains the well-established observation (Huang et al., 2021; Park et al., 2023; Sun et al., 2022; Tack et al., 2020) that OOD features tend to reside closer to the origin.
@@
Notably, prior methods, e.g., KNN (Sun et al., 2022), focus on ID clustering but do not explicitly consider feature distance to the origin. Such approaches fall short in scenarios like ImageNet benchmarks but yield superior performance in CIFAR-10 benchmarks in Table 1(a). Conversely, methods such as Energy (Liu et al., 2020), Energy-based ASH (Djurisic et al., 2022), and, Energy-based Scale (Xu et al., 2023) inherently utilize feature distance to the origin by considering log-sum-exp of logits, yet largely overlook ID clustering. These approaches excel in some scenarios, e.g., ImageNet, but perform sub-optimally in others, e.g., CIFAR-10 (see Table 1(a)). Through the lens of Neural Collapse, we explain, connect, and complete prior methods under a holistic view, leading to improved efficiency and generalizability validated across diverse experiments.
@@
  • Experimental Analysis: We evaluate $\mathtt{NCI}$ across diverse classification tasks – CIFAR-10, CIFAR-100, ImageNet – and model architectures – ResNet, DenseNet (Huang et al., 2017), ViT (Dosovitskiy et al., 2020), Swin (Liu et al., 2022). Our $\mathtt{NCI}$ practically matches the latency of vanilla softmax-confidence detector while largely mitigating the performance discrepancy of existing methods across different classification tasks and model architectures.
@@ 3 OOD Detection through the Lens of Neural Collapse
In this section, we re-examine the observation in Lee et al. (2018) and Sun et al. (2022) that ID features tend to form clusters while OOD features deviate from the clusters. We understand the phenomenon as a precursor to the Neural Collapse (Papyan et al., 2020) convergence. And leveraging the landscape revealed in Neural Collapse, we examine:
@@ 3.1 Neural Collapse: Convergence of Training Features
We first remark on (NC2) that an ETF achieves the maximum separation possible for globally centered equiangular vectors (Papyan et al., 2020) and extends in space, as visualized in Figure 1 Right. Since training features converge towards an ETF, they need to have sufficient norms to accommodate the spatial arrangement.
@@ 3.2 Geometric Structure of the Clusters of ID Features
Additionally, we combine our observation with the results from Zhu et al. (2021), showing that the weight vectors form an ETF during training. Our observed proximity to the weight vectors thus also validates the clustering of ID features near an ETF and the divergence of OOD from this structure. The lack of structure and incentives to extend in space explains the relatively smaller norm of OOD features.
@@ 3.3 Out-of-Distribution Detection
While $\mathtt{pScore}$ enhances efficiency and effectiveness, its performance is intrinsically contingent on the strength of ID clustering. Such contingency, widely exhibited by clustering-based methods (Lee et al., 2018; Sun et al., 2022; Tack et al., 2020), poses challenges on classifiers with less pronounced ID clustering, such as ImageNet ResNet-50 in Section 4.1. To mitigate such discrepancy, we complement $\mathtt{pScore}$ by considering the distance of ID clusters to the origin. Specifically, we enhance our proximity score by incorporating feature norms to filter out OOD near the origin, as illustrated in Figure 1 Right. Taking $\mathtt{L1}$ norm as an example, we define our detection score as $\mathtt{pScore}+\alpha\|\bm{h}\|_{1}$, where $\alpha$ controls the filtering strength and can be selected from a validation set as detailed in Section 4. We refer readers to Section 4.3 for the effect of different orders of $p$-norm. Thresholding on the detection score, we have Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$): A lower score indicates a higher chance of OOD-ness.
@@ 4 Experiments
(a) $\mathtt{NCI}$ remains effective while baselines exhibit discrepancy. $\uparrow$ and $\downarrow$ denotes better performance. Bold indicates the best results while underline denotes the 2nd and 3rd best results. Methods with * are hyperparameter-free. Scores, except for the most recent baselines – $\mathtt{fDBD}$, $\mathtt{NECO}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – are copied from OpenOOD report (Zhang et al., 2023).
@@
In this section, we extensively demonstrate the versatility and efficiency of $\mathtt{NCI}$ across classification tasks: CIFAR-10, CIFAR-100 (see App. D), ImageNet, as well as model architectures: ResNet, DenseNet (see App. D), ViT, Swin. We compare $\mathtt{NCI}$ against thirteen baseline methods, including the most recent ones, and demonstrate that $\mathtt{NCI}$ effectively mitigates the generalization discrepancies of existing methods. Following the OpenOOD benchmark (Zhang et al., 2023), we evaluate on six OOD sets for CIFAR-10 and CIFAR-100 classifiers and five for ImageNet classifiers. Performance is evaluated using two widely recognized metrics: the False Positive Rate at 95% True Positive Rate (FPR95) and the Area Under the Receiver Operating Characteristic Curve (AUROC). Lower FPR95 and higher AUROC values indicate better performance. We also report the per-image inference latency (in milliseconds) evaluated on a Tesla T4 GPU. In our experiments, other than the ablation study in Section 4.3, we use the $L1$-norm as the filtering term and select the filtering strength $\alpha$ from $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ based on a validation set generated per pixel from Gaussian $N(0,1)$, following Sun et al. (2021) and Sun and Li (2022). In Section 4.3, we explore different choices of norm order and examine the effect of filtering on alternative clustering-based method. For detailed setups, please see Appendix A. We emphasize that our method and all baselines are post-hoc methods, while all models used are off-the-shelf and do not undergo prolonged training.
@@ 4.1 Versatility across Classification Tasks
We first assess the performance of $\mathtt{NCI}$ and baselines across CIFAR-10 and ImageNet classification tasks. The two tasks provide an ideal test bed for evaluating versatility, as they drastically differ in input resolution, number of classes, and classification accuracy. We use the ResNets from the OpenOOD benchmark (Zhang et al., 2023): the CIFAR-10 classifier is a ResNet-18 with an accuracy of 95.06% and the ImageNet classifier is a ResNet-50 with an accuracy of 76.65%. Based on validation results, we set the filter strength $\alpha$ of the $L1$-norm to $10^{-2}$ for CIFAR-10 experiments and $10^{-3}$ for ImageNet experiments.
Datasets For CIFAR-10 experiments, We follow the OpenOOD split of CIFAR-10 test set and evaluate on the OpenOOD benchmarks, including CIFAR-100 (Krizhevsky et al., 2009), Tiny ImageNet (Le and Yang, 2015), MNIST (Deng, 2012), SVHN (Netzer et al., 2011), Texture (Cimpoi et al., 2014), and Places365 (Zhou et al., 2017). For ImageNet experiments, we follow the OpenOOD split of ImageNet test set and evaluate on the OpenOOD benchmarks, including SSB-hard (Vaze et al., 2021), NINCO (Bitterwolf et al., 2023), iNaturalist (Van Horn et al., 2018), Texture (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022).
Baselines In Table 1(a), we compare our method with thirteen baselines. Some baselines focus more on the CIFAR-10 Benchmark while others focus more focused on the Imagenet Benchmark. Based on performance, we categorize the baselines, besides the vanilla confidence-based $\mathtt{MSP}$ (Hendrycks and Gimpel, 2016), into two groups: the “CIFAR-10 Strong" baselines, including $\mathtt{ODIN}$ (Liang et al., 2018), $\mathtt{Energy}$ (Liu et al., 2020), $\mathtt{Mahalanobis}$ (Lee et al., 2018), $\mathtt{KNN}$ (Sun et al., 2022), $\mathtt{ViM}$ (Wang et al., 2022), and $\mathtt{fDBD}$ (Liu and Qin, 2023); the “ImageNet Strong" baselines, including $\mathtt{GradNorm}$ (Huang et al., 2021), $\mathtt{NECO}$ (Ammar et al., 2023), $\mathtt{React}$ (Sun et al., 2021), $\mathtt{Dice}$ (Sun and Li, 2022), $\mathtt{ASH}$ (Djurisic et al., 2022), $\mathtt{Scale}$ (Xu et al., 2023). For details of the baselines, please see Appendix C.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{NECO}$: Like our $\mathtt{NCI}$, $\mathtt{NECO}$ (Ammar et al., 2023) is motivated by Neural Collapse. Similar to our $\mathtt{NCI}$ with filtering, $\mathtt{NECO}$ incorporates distance to the origin using max-logit. However, $\mathtt{NECO}$, exclusively focuses on the subspace analysis of features and does not utilize the classification head. Such subspace operations require expensive matrix multiplication, resulting in a noticeable inference latency shown in Table 1(b). Conversely, our $\mathtt{NCI}$ explores the interplay between features and the classification head, and thus integrates class-specific information revealed by Neural Collapse. Compared to $\mathtt{NECO}$, our $\mathtt{NCI}$ improves in both efficiency and effectiveness.
@@ 4.2 Versatility across Model Architectures
Now we assess the performance of $\mathtt{NCI}$ and baselines across different architectures. Particularly, we study two transformer-based models: ViT B/16 (Dosovitskiy et al., 2020) and Swin-v2 (Liu et al., 2022). Both transformers are finetuned on ImageNet, achieving an accuracy of 81.14% and 82.94% respectively. We follow the setup of the OpenOOD ImageNet Benchmark, detailed in Section 4.1. And based on validation results, we set the filter strength $\alpha$ of the $L1$ norm to $10^{-3}$ for both classifiers. In Table 2, we observe strong baselines on ViT – $\mathtt{KNN}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – exhibit discrepancy when applied to Swin v2, echoing the observations in Ammar et al. (2023). Conversely, our $\mathtt{NCI}$, even without filtering, improves baseline performance on Swin v2 while maintaining superior effectiveness on ViT. The addition of filtering further enhances overall generalizability. Our experiments on ResNet (Section 4.1) and DenseNet (Appendix D) further validate the versatility of $\mathtt{NCI}$ across model architectures.